OpenStudy (anonymous):

solve the following system by substitution. substitute the expression in the second equation into the first equation and solve. 2x-4y= -22 x= -2y+5

OpenStudy (anonymous):

2x-4y=-22 x=-2y+5 substitute second equation in first 2(-2y+5)=-22 multiplying -4y+10=-22 -4y=-32 y=8 put the value of y in any of the above two equations 2x-4(8)=-22 2x-32=-22 2x=10 x=5 x=5,y=8